A metal surface is illuminated by light with a wavelength of 350 nm . The maximum kinetic energy of the emitted electrons is found to be 1.90 eV .

What is the maximum electron kinetic energy if the same metal is illuminated by light with a wavelength of 250 nm ?
Answer #1

so we are given

L1=350nm

E1=1.90eV

E2=?

L2=250nm

max. K.E of electron = energy of incident photon - work function

E=hc/L-W

for

E1=hc/L1-W

1.90=hc/350-W

W=hc/350-1.90

E2=hc/L2-W

E2=hc/250-W

E2=hc/250-(hc/350-1.90)

=hc/250-hc/350+1.90

=0.0011hc+1.90

hc = 1240 eV nm

E2=3.26eV---answer